Supplemental Exercises 1. The densities of magnesium and palladium are 1740 kg/m3 and 11900 kg/m respectively [This data was taken from www.webelemen om.] Convert these densities to units of g/mL. 2. type of metal you used in the experiment. If the cylinder used in this experiment had been made of magnesium, but had the same mass, would it have been bigger smaller? If cylinder used in this experiment had been made of but had the same mass, would it have been bigger or smaller? Be sure to state which metal you used.
Please solve both questions

For question #2 the metal we used was brass and it weighed 8.6g/mL
Answer #1

We know 1 kg =1000 gm

1 m^3=1000L=10^6 ml

So...1740kg/m^3=1740*10^-3 gm/ml

And for palladium its 11900*10^-3gm/ml